OpenStudy (anonymous):

Identify the roots and state the multiplicity x^6 - 12x^4 + 48x^2 -64 = 0

OpenStudy (anonymous):

Thank God! I finally solved it. x^6 -12x^4+48x^2-64=0 (x^2-4) (x^2-4) (x^2-4)=0 x^2-4=0 x^2=4 x+-2 or x=2 with the multiplicity of 3. It was extremely hard. Hope I helped.

OpenStudy (anonymous):

Oh, before writing (x^2-4(x^2-4) (x^2-4)=0 you should put (x^2-4)^3